# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v1

@@ Abstract
We explore the thermodynamics of quantum processes (quantum channels) by axiomatically introducing the free energy for channels, defined via the quantum relative entropy with an absolutely thermal channel whose fixed output is in equilibrium with a thermal reservoir. This definition finds strong support through its operational interpretations in designated quantum information and thermodynamic tasks. We construct a resource theory of athermality for quantum processes, where free operations are Gibbs preserving superchannels and golden units are unitary channels with respect to absolutely thermal channel having fully degenerate output Hamiltonian. We exactly characterize the one-shot distillation and formation of quantum channels using hypothesis-testing and max-relative entropy with respect to the absolutely thermal channel. These rates converge asymptotically to the channel free energy (up to a multiplicative factor of half the inverse temperature), establishing its operational meaning and proving the asymptotic reversibility of the athermality. We show the direct relation between the resource theory of athermality and quantum information tasks such as private randomness and purity distillation and thermodynamic tasks of erasure and work extraction. Our work connects the core thermodynamic concepts of free energy, energy, entropy, and maximal extractable work of quantum processes to their information processing capabilities.
@@ I Introduction
Quantum computation and information processing are fundamentally built upon quantum processes [1, 2, 3, 4, 5], which are the physical transformations and evolutions that quantum systems undergo. Studying the thermodynamic aspects of these systems and processes through an information-theoretic approach offers essential insights into the fundamental limitations on preparation, manipulation, and measurement of quantum resources, especially when faced with energetic constraints [6, 7, 8, 9, 10, 11]. Free energy stands as a core concept in thermodynamics, and its extension to quantum states provides insight to limitations on allowed transformations of states [12, 13, 14, 15, 16] and possible utilization for thermodynamic engines and circuits [17, 18, 19, 20].
Figure 1: A bipartite system with a non-interacting Hamiltonian $\widehat{H}_{R}+\widehat{H}_{A}$ is in state $\psi_{RA^{\prime}}$, in contact with a bath at inverse temperature $\beta$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ acts locally on $A^{\prime}$ such that $\widehat{H}^{\rm int}_{RA}=0$. To extract work, we partially thermalize the channel output, $\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})\to\psi_{R}\otimes\gamma_{A}$. We consider partial thermalization composed of three steps: (1) Decoupling: $\mathcal{N}(\psi_{RA^{\prime}})=:\rho_{RA}\to\rho_{R}\otimes\rho_{A}$, (2) Quenching: The Hamiltonian is instantaneously changed from $\widehat{H}_{A}$ to $\beta^{-1}\ln\rho_{A}$ such that $\rho_{A}$ is an equilibrium state. (3) Isothermal quasistatic transformation of the Hamiltonian back to $\widehat{H}_{A}$. The extractable work from each step of the process is written in the blue color. The thermodynamic utility of the channel is quantified as the maximal extractable work via the partial thermalization process, see Theorem 6.
@@
Thermodynamic free energy $F_{c}^{\beta}(\rho)=\beta^{-1}D(\rho\|\widehat{\gamma}^{\beta})$ $F_{c}^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$
Resource-theoretic free energy $F^{\beta}(\rho)=F^{\beta}_{c}(\rho)-F^{\beta}_{c}(\gamma^{\beta})=\beta^{-1}D(\rho\|{\gamma}^{\beta})$ $F^{\beta}[\mathcal{N}]=F_{c}^{\beta}[\mathcal{N}]-F_{c}^{\beta}[\mathcal{T}^{\beta}]=\beta^{-1}D[\mathcal{N}\|{\mathcal{T}}^{\beta}]$
@@
Maximal extractable work $W^{\rm ext}_{\rho\to\gamma}=F^{\beta}_{c}(\rho)-F_{c}^{\beta}(\gamma^{\beta})=F^{\beta}(\rho)$ $W^{\rm ext}_{\mathcal{N}\to\mathcal{T}^{\beta}}=F_{c}^{\beta}[\mathcal{N}]-F_{c}^{\beta}[\mathcal{T}^{\beta}]=F^{\beta}[\mathcal{N}]$
Helmholtz equation $F^{\beta}_{c}(\rho)=E(\rho)-\beta^{-1}S(\rho)$ $F^{\beta}_{c}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}]$
@@
Table 1: We summarize and compare mathematical expressions for elementary thermodynamic concepts for quantum states vs channels. We consider thermal reservoir to be at inverse temperature $\beta$ and $\widehat{H}$ denotes the Hamiltonian of the designated system. $\widehat{\gamma}^{\beta}$ is a thermal (Gibbs) operator and $\gamma^{\beta}$ is a thermal state. $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\omega_{A}$, $\mathcal{T}^{\beta}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\gamma^{\beta}_{A}$, $\widehat{\mathcal{T}}^{\beta}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\widehat{\gamma}^{\beta}$. $D(\rho\|\sigma):=\lim_{\varepsilon\to 0^{+}}\operatorname{tr}[\rho(\ln\rho-\ln(\sigma+\varepsilon\mathbbm{1}))]$ is (Umegaki) relative entropy between states $\rho,\sigma$ and $D[\mathcal{N}\|\mathcal{M}]:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}})\|\operatorname{id}_{R}\otimes\mathcal{M}(\psi_{RA^{\prime}}))$ is relative entropy between channels $\mathcal{N}_{A^{\prime}\to A},\mathcal{M}_{A^{\prime}\to A}$. We assume $\rho$ is a density operator of system $A^{\prime}$ and $\mathcal{N}$ is a quantum channel from $A^{\prime}\to A$. The content for quantum states are known in literature [12, 21, 22, 23], prior to our current work. The expression of the entropy of quantum channels is from [24, 25]. See Section I.2 for the outline of the paper.
@@ I.1 Main Results
The concept of the free energy is fundamental and instrumental in thermodynamics. We extend this notion to quantum processes which describe physical transformations of quantum systems in quantum theory. We leverage axiomatic and information-theoretic approaches to define meaningful free energy functions of quantum channels (Theorem 1) that find operational interpretations in quantum information processing and thermodynamical tasks. We establish relations between the elementary thermodynamic quantities like free energy, energy, entropy, and operational quantity like extractable work to shed light on the thermodynamics of quantum processes (see Table 1).
@@
We define the resource-theoretic generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ as the appropriate generalized channel divergence functions, $\mathbf{F}^{\beta}[\mathcal{N}]:=\beta^{-1}\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ [25, 30], between the channel $\mathcal{N}$ and an absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$. The channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ thermalizes all its inputs such that the output state is a thermal state $\gamma^{\beta}_{A}$ in equilibrium with the bath at an inverse temperature $\beta>0$, $\mathcal{T}^{\beta}(\cdot)=\operatorname{tr}[\cdot]\gamma^{\beta}$. Some appropriate generalized channel divergences for our purpose are shown to be the Umegaki relative entropy and max-relative entropy. We also discuss smoothed versions of free energy functions derived from these channel divergences and the $\varepsilon$-hypothesis-testing relative entropy.
@@
We develop the dynamical resource theory of athermality for (square) quantum channels. The absolutely thermal channel $\mathcal{T}^{\beta}$ is considered a free object, where $\mathcal{T}^{\beta}$ depends on the inverse temperature of the bath and the Hamiltonian of the output system under consideration. This is a natural assumption because of the inherent thermalization process that equilibrates quantum systems with the bath. Even near isolation of quantum systems from the bath for short time requires engineering and interventions, and perfect closed quantum dynamics are idealistic. The free operations are Gibbs preserving superchannels $\Theta^{\beta}$ which map absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ to absolutely thermal channel $\mathcal{T}^{\beta}_{B^{\prime}\to B}$. A quantum channel $\mathcal{N}$ is athermal if it isn’t asbolutely thermal $\mathcal{T}^{\beta}$ and is deemed resourceful. For clarity, we use $(\mathcal{N},\mathcal{T}^{\beta})$ to discuss the thermodynamic resourcefulness of $\mathcal{N}$ with respect to $\mathcal{T}^{\beta}$ [31].
@@
The resource-theoretic (and thermodynamic) free energy and max-free energy are maximum for a square quantum channel if and only if it’s a unitary channel (Theorem 2). Among all possible Hamiltonians, the free energy of a unitary channel $\mathcal{U}_{A^{\prime}\to A}$ is least if the output Hamiltonian is fully degenerate, $\widehat{H}_{A}=c\mathbbm{1}_{A}$ for $c\in\mathbbm{R}$. With certain physical motivation, we advocate that a golden unit of the dynamical resource theory of athermality is the resource channel $(\operatorname{id}_{m},\mathcal{R}^{\gamma})$, where ${\mathcal{N}}_{m}$ denotes that input and output Hilbert spaces of channel $\mathcal{N}$ are $m$-dimensional and $\mathcal{R}^{\pi}$ is the uniformly mixing channel that only outputs the maximally mixed state $\pi$. All unitary resource channels $(\mathcal{U}_{m},\mathcal{R}^{\pi})$ are equivalent (equally resourceful) to $(\operatorname{id}_{m},\mathcal{R}^{\pi})$. In fact, the interconversion between $(\mathcal{U},\mathcal{T}^{\beta})\leftrightarrow(\mathcal{V},\mathcal{T}^{\beta})$ is possible via Gibbs preserving superchannels (free operations) for any pair of unitary channels $\mathcal{U}$ and $\mathcal{V}$.
@@
We find that the single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, up to allowed error $\varepsilon\in(0,1)$ under the action of Gibbs preserving superchannels, are $\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$, respectively (Theorem 3). We have $F^{\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$, which provide exact operational meaning to the resource-theoretic hypothesis-testing free energy and the max-free energy of a quantum channel. The asymptotic athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, irrespective of adaptive or nonadaptive strategies, are equal to $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ (Theorem 4). The resource-theoretic free energy of a channel $\mathcal{N}$, $F^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, and thus its asymptotic athermality distillation and formation rates are both equal to $\frac{\beta}{2}F^{\beta}[\mathcal{N}]$, providing the exact operational interpretation to the resource-theoretic free energy of a quantum channel. Both asymptotic athermality distillation and formation rates of a quantum channel are the same, implying that the athermality of quantum channels under Gibbs preserving superchannels is (asymptotically) reversible.
@@
We get the resource theory of purity for quantum channels as a special case in the limit $\beta\to 0^{+}$ or Hamiltonian of the channel output is trivial. We discuss the dependence of the free energy of a quantum channel on quantum transmission ability of the channel and free energy of its output (Proposition 5). We elucidate on the novel relations between the free energy of a quantum channel and its energy and entropy. The thermodynamic free energy $F_{c}[\mathcal{N}]:=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$, where $\widehat{\mathcal{T}}^{\beta}$ is a replacer (or preparation) map that always outputs thermal operator $\widehat{\gamma}^{\beta}$ (an unnormalized thermal state), of a quantum channel $\mathcal{N}$ and its entropy $S[\mathcal{N}]=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$ [24, 25], where $\mathcal{R}^{\mathbbm{1}}(\cdot)=\operatorname{tr}[\cdot]\mathbbm{1}$, and thermodynamic energy $E[\mathcal{N}]$ satisfy the relation $F_{c}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}]$ when the channel output is noninteracting with its reference (Theorem 5).
@@
We prove that the resource-theoretic free energy of a quantum channel $\mathcal{N}$ is the difference of the thermodynamic free energy of the channel and the absolutely thermal channel, $F[\mathcal{N}]=F_{c}[\mathcal{N}]-F_{c}[\mathcal{T}^{\beta}]$, and has operational interpretation as the optimal (maximum) thermodynamic utility of the channel for maximal extractable work (Theorem 6, see Fig. 1). We believe our framework, results, and discussions will provide deeper insights on thermodynamical aspects of quantum processes and interconnectedness with their quantum information processing capabilities.

### v2

@@ Abstract
We explore the thermodynamics of quantum processes (quantum channels) by axiomatically introducing the free energy for channels, defined via the quantum relative entropy with an absolutely thermal channel whose fixed output is in equilibrium with a thermal reservoir. This definition finds strong support through its operational interpretations in designated quantum information and thermodynamic tasks. We construct a resource theory of athermality for quantum processes, where free operations are Gibbs preserving superchannels and golden units are unitary channels with respect to absolutely thermal channel having fully degenerate output Hamiltonian. We exactly characterize the one-shot distillation and formation of quantum channels using hypothesis-testing and max-relative entropy with respect to the absolutely thermal channel. These rates converge asymptotically to the channel free energy (up to a multiplicative factor of half the inverse temperature), establishing its operational meaning and proving the asymptotic reversibility of the athermality. We show the direct relation between the resource theory of athermality and quantum information tasks such as private randomness and purity distillation, and thermodynamic tasks of erasure and work extraction. Our work connects the core thermodynamic concepts of free energy, energy, entropy, and maximal extractable work of quantum processes to their information processing capabilities.
@@ I Introduction
Quantum computation and information processing are fundamentally built upon quantum processes [1, 2, 3, 4, 5], which are the physical transformations and evolutions that quantum systems undergo. Studying the thermodynamic aspects of these systems and processes through an information-theoretic approach offers essential insights into the fundamental limitations on preparation, manipulation, and measurement of quantum resources, especially when faced with energetic constraints [6, 7, 8, 9, 10, 11]. Free energy stands as a core concept in thermodynamics, and its extension to quantum states provides insight into limitations on allowed transformations of states [12, 13, 14, 15, 16] and possible utilization for thermodynamic engines and circuits [17, 18, 19, 20].
Figure 1: A bipartite system with a non-interacting Hamiltonian $\widehat{H}_{R}+\widehat{H}_{A}$ is in state $\psi_{RA^{\prime}}$, in contact with a bath at inverse temperature $\beta$. A quantum channel $\mathcal{N}_{A^{\prime}\to A}$ acts locally on $A^{\prime}$ such that $\widehat{H}^{\rm int}_{RA}=0$. To extract work, we partially thermalize the channel output, $\mathcal{N}_{A^{\prime}\to A}(\psi_{RA^{\prime}})\to\psi_{R}\otimes\gamma_{A}$. We consider partial thermalization composed of three steps: (1) Decoupling: $\mathcal{N}(\psi_{RA^{\prime}})=:\rho_{RA}\to\rho_{R}\otimes\rho_{A}$, (2) Quenching: The Hamiltonian is instantaneously changed from $\widehat{H}_{A}$ to $\beta^{-1}\ln\rho_{A}$ such that $\rho_{A}$ is an equilibrium state. (3) Isothermal quasistatic transformation of the Hamiltonian back to $\widehat{H}_{A}$. The extractable work from each step of the process is written in blue color. The thermodynamic utility of the channel is quantified as the maximal extractable work via the partial thermalization process, see Theorem 5.
@@
Thermal free energy $F_{\rm T}^{\beta}(\rho)=\beta^{-1}D(\rho\|\widehat{\gamma}^{\beta})$ $F_{\rm T}^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$
Resource-theoretic free energy $F^{\beta}(\rho)=F^{\beta}_{\rm T}(\rho)-F^{\beta}_{\rm T}(\gamma^{\beta})=\beta^{-1}D(\rho\|{\gamma}^{\beta})$ $F^{\beta}[\mathcal{N}]=F_{\rm T}^{\beta}[\mathcal{N}]-F_{\rm T}^{\beta}[\mathcal{T}^{\beta}]=\beta^{-1}D[\mathcal{N}\|{\mathcal{T}}^{\beta}]$
@@
Maximal extractable work $W^{\rm ext}_{\rho\to\gamma}=F^{\beta}_{\rm T}(\rho)-F_{\rm T}^{\beta}(\gamma^{\beta})=F^{\beta}(\rho)$ $W^{\rm ext}_{\mathcal{N}\to\mathcal{T}^{\beta}}=F_{\rm T}^{\beta}[\mathcal{N}]-F_{\rm T}^{\beta}[\mathcal{T}^{\beta}]=F^{\beta}[\mathcal{N}]$
Helmholtz equation $F^{\beta}_{\rm T}(\rho)=E(\rho)-\beta^{-1}S(\rho)$ $F^{\beta}_{\rm T}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}]$
@@
Table 1: We summarize and compare mathematical expressions for elementary thermodynamic concepts for quantum states vs channels. We consider the thermal reservoir to be at inverse temperature $\beta$ and $\widehat{H}$ denotes the Hamiltonian of the designated system. $\widehat{\gamma}^{\beta}$ is a thermal (Gibbs) operator and $\gamma^{\beta}$ is a thermal state. $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\omega_{A}$, $\mathcal{T}^{\beta}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\gamma^{\beta}_{A}$, $\widehat{\mathcal{T}}^{\beta}_{A^{\prime}\to A}(\cdot):=\operatorname{tr}[\cdot]\widehat{\gamma}^{\beta}$. $D(\rho\|\sigma):=\lim_{\varepsilon\to 0^{+}}\operatorname{tr}[\rho(\ln\rho-\ln(\sigma+\varepsilon\mathbbm{1}))]$ is (Umegaki) relative entropy between states $\rho,\sigma$ and $D[\mathcal{N}\|\mathcal{M}]:=\sup_{\psi\in\operatorname{St}(RA^{\prime})}D(\operatorname{id}_{R}\otimes\mathcal{N}(\psi_{RA^{\prime}})\|\operatorname{id}_{R}\otimes\mathcal{M}(\psi_{RA^{\prime}}))$ is relative entropy between channels $\mathcal{N}_{A^{\prime}\to A},\mathcal{M}_{A^{\prime}\to A}$. We assume $\rho$ is a density operator of system $A^{\prime}$ and $\mathcal{N}$ is a quantum channel from $A^{\prime}\to A$. The content for quantum states are known in literature [12, 21, 22, 23], prior to our current work. The expression of the entropy of quantum channels is from [24, 25]. See Section I.2 for the outline of the paper.
@@
Note added. This paper is companion to “Thermodynamic work capacity of quantum information processing” [47].
@@ I.1 Main Results
The concept of free energy is fundamental and instrumental in thermodynamics. We extend this notion to quantum processes which describe physical transformations of quantum systems in quantum theory. We leverage axiomatic and information-theoretic approaches to define meaningful free energy functions of quantum channels (Theorem 1) that find operational interpretations in quantum information processing and thermodynamical tasks. We establish relations between the elementary thermodynamic quantities like free energy, energy, entropy, and operational quantities like extractable work to shed light on the thermodynamics of quantum processes (see Table 1).
@@
We define the resource-theoretic generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ as the inverse temperature times the appropriate generalized channel divergence functions, $\mathbf{F}^{\beta}[\mathcal{N}]:=\beta^{-1}\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ [25, 30], between the channel $\mathcal{N}$ and an absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$. The channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ thermalizes all its inputs such that the output state is a thermal state $\gamma^{\beta}_{A}$ in equilibrium with the bath at an inverse temperature $\beta>0$, $\mathcal{T}^{\beta}(\cdot)=\operatorname{tr}[\cdot]\gamma^{\beta}$. Some appropriate generalized channel divergences for our purpose are shown to be the Umegaki relative entropy and max-relative entropy. We also discuss smoothed versions of free energy functions derived from these channel divergences and the $\varepsilon$-hypothesis-testing relative entropy.
@@
We develop the dynamical resource theory of athermality for (square) quantum channels. The absolutely thermal channel $\mathcal{T}^{\beta}$ is considered a free object, where $\mathcal{T}^{\beta}$ depends on the inverse temperature of the bath and the Hamiltonian of the output system under consideration. This is a natural assumption because of the inherent thermalization process that equilibrates quantum systems with the bath. Even near isolation of quantum systems from the bath for a short time requires engineering and interventions, and perfect closed quantum dynamics are idealistic. The free operations are Gibbs preserving superchannels $\Theta^{\beta}$ which map absolutely thermal channel $\mathcal{T}^{\beta}_{A^{\prime}\to A}$ to absolutely thermal channel $\mathcal{T}^{\beta}_{B^{\prime}\to B}$ of the same inverse temperature $\beta$. A quantum channel $\mathcal{N}$ is athermal if it is not absolutely thermal $\mathcal{T}^{\beta}$ and is deemed resourceful. For clarity, we use $(\mathcal{N},\mathcal{T}^{\beta})$ to discuss the thermodynamic resourcefulness of $\mathcal{N}$ with respect to $\mathcal{T}^{\beta}$ [31].
@@
The resource-theoretic (and thermodynamic) free energy and max-free energy are maximum for a square quantum channel if and only if it’s a unitary channel (Proposition 2). Among all possible Hamiltonians, the free energy of a unitary channel $\mathcal{U}_{A^{\prime}\to A}$ is least if the output Hamiltonian is fully degenerate, $\widehat{H}_{A}=c\mathbbm{1}_{A}$ for $c\in\mathbbm{R}$. As the Hamiltonian can be arbitrary (and large), we advocate that a golden unit of the dynamical resource theory of athermality is the resource channel $(\operatorname{id}_{m},\mathcal{R}^{\pi})$, where ${\mathcal{N}}_{m}$ denotes that input and output Hilbert spaces of channel $\mathcal{N}$ are $m$-dimensional and $\mathcal{R}^{\pi}$ is the uniformly mixing channel that only outputs the maximally mixed state $\pi$. All unitary resource channels $(\mathcal{U}_{m},\mathcal{R}^{\pi})$ are equivalent (equally resourceful) to $(\operatorname{id}_{m},\mathcal{R}^{\pi})$. In fact, the interconversion between $(\mathcal{U},\mathcal{T}^{\beta})\leftrightarrow(\mathcal{V},\mathcal{T}^{\beta})$ is possible via Gibbs preserving superchannels (free operations) for any pair of unitary channels $\mathcal{U}$ and $\mathcal{V}$.
@@
We find that the single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, up to allowed error $\varepsilon\in(0,1)$ under the action of Gibbs preserving superchannels, are $\frac{1}{2}D_{\infty}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$ and $\frac{1}{2}D_{H}^{\varepsilon}[\mathcal{N}\|\mathcal{T}^{\beta}]$, respectively (Theorem 2). We have $F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$, which provide exact operational meaning to the resource-theoretic hypothesis-testing free energy and the max-free energy of a quantum channel. The asymptotic athermality distillation and formation rates of a quantum resource channel $(\mathcal{N},\mathcal{T}^{\beta})$, under parallel uses of channels, are equal to $\frac{1}{2}D[\mathcal{N}\|\mathcal{T}^{\beta}]$ (Theorem 3). The resource-theoretic free energy of a channel $\mathcal{N}$, $F^{\beta}[\mathcal{N}]=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, and thus its asymptotic athermality distillation and formation rates are both equal to $\frac{\beta}{2}F^{\beta}[\mathcal{N}]$, providing the exact operational interpretation to the resource-theoretic free energy of a quantum channel. Both asymptotic athermality distillation and formation rates of a quantum channel are the same, implying that the athermality of quantum channels under Gibbs preserving superchannels is (asymptotically) reversible.
@@
We get the resource theory of purity for quantum channels as a special case in the limit $\beta\to 0^{+}$ or Hamiltonian of the channel output is trivial. We discuss the dependence of the free energy of a quantum channel on the quantum transmission ability of the channel and free energy of its output (Proposition 4). We elucidate on the novel relations between the free energy of a quantum channel and its energy and entropy. The thermal free energy $F^{\beta}_{\rm T}[\mathcal{N}]:=\beta^{-1}D[\mathcal{N}\|\widehat{\mathcal{T}}^{\beta}]$, where $\widehat{\mathcal{T}}^{\beta}$ is a replacer (or preparation) map that always outputs thermal operator $\widehat{\gamma}^{\beta}$ (an unnormalized thermal state), of a quantum channel $\mathcal{N}$ and its entropy $S[\mathcal{N}]=-D[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$ [24, 25], where $\mathcal{R}^{\mathbbm{1}}(\cdot)=\operatorname{tr}[\cdot]\mathbbm{1}$, and thermodynamic energy $E[\mathcal{N}]$ satisfy the relation $F^{\beta}_{\rm T}[\mathcal{N}]\leq E[\mathcal{N}]-\beta^{-1}S[\mathcal{N}]$ when the channel output is noninteracting with its reference (Theorem 4).
@@
We prove that the resource-theoretic free energy of a quantum channel $\mathcal{N}$ is the difference of the thermal free energy of the channel and the absolutely thermal channel, $F^{\beta}[\mathcal{N}]=F^{\beta}_{\rm T}[\mathcal{N}]-F^{\beta}_{\rm T}[\mathcal{T}^{\beta}]$, and has operational interpretation as the optimal (maximum) thermodynamic utility of the channel for maximal extractable work (Theorem 5, see Fig. 1). We believe our framework, results, and discussions will provide deeper insights into the thermodynamical aspects of quantum processes and interconnectedness with their quantum information processing capabilities.